OpenStudy (anonymous):

Solve: 5x -2 = 2x + 7

geerky42 (geerky42):

Subtract both sides by 2x: 5x - 2x - 2 = 2x - 2x + 7 3x - 2 = 7 Add both sides by 2: 3x - 2 + 2 = 7 + 2 3x = 9 Divide both sides by 3: 3x/3 = 9/3 \(\boxed{x = 3}\)

OpenStudy (anonymous):

Awesome! Thanks I'm new to this site and already by your help I'm loving it! Thanks for your help I really appreciate it. :)

geerky42 (geerky42):

You're welcome :)

OpenStudy (anonymous):

5x-2=2x+7 Subtract 2x from 5x which makes 2x cancel out Then now you have 3x Drop down the rest of the equation which is 3x-2=7 Add 2 to both sides This makes 2 cancel out 7+2=9 drop down the rest of the equation which is 3x=9 Divide 3 from both sides you should have x=3 Hope I Helped